# General Perturbative Expansion and Relations with Unitarity Triangles

### v1

@@ III Summary
In this letter, using a rephasing invariant formula $\delta=\arg[{V_{ud}V_{us}V_{cb}V_{tb}/V_{ub}\det V_{\rm CKM}}]$, we evaluate the CP phase $\delta$ of the CKM matrix $V_{\rm CKM}$ perturbatively for small quark mixing angles $s_{ij}^{u,d}$ with associated phases $\rho_{ij}^{u,d}$. Consequently, we derived a relation $\delta\simeq\arg[\Delta s_{12}\Delta s_{23}/(\Delta s_{13}-s^{u}_{12}e^{-i\rho^{u}_{12}}\Delta s_{23})]$ with $\Delta s_{ij}\equiv s^{d}_{ij}e^{-i\rho^{d}_{ij}}-s^{u}_{ij}e^{-i\rho^{u}_{ij}}$. Such a result represents the analytic behavior of the CKM phase. The uncertainty in the relation is of order $O(\lambda^{2})\sim 4\%$, which is comparable to the current experimental precision. Comparisons with experimental data suggest that the hypothesis of some CP phases being maximal.
We also discussed relationships between the phase $\delta$ and unitarity triangles. As a result, several relations between the angles $\alpha,\beta,\gamma$ and $\delta$ are identified through other invariants $V_{il}V_{jm}V_{kn}/\det V_{\rm CKM}$. These general perturbative relations broadly cover phenomenological calculations, and therefore, the presented results have wide applicability in studies of flavor physics and CP violation.

### v2

@@ III Summary
In this letter, using a rephasing invariant formula $\delta=\arg[{V_{ud}V_{us}V_{cb}V_{tb}/V_{ub}\det V_{\rm CKM}}]$, we evaluate the CP phase $\delta$ in the CKM matrix $V_{\rm CKM}$ perturbatively for small quark mixing angles $s_{ij}^{u,d}$ with associated phases $\rho_{ij}^{u,d}$. Consequently, we derived a relation $\delta\simeq\arg[\Delta s_{12}\Delta s_{23}/(\Delta s_{13}-s^{u}_{12}e^{-i\rho^{u}_{12}}\Delta s_{23})]$ with $\Delta s_{ij}\equiv s^{d}_{ij}e^{-i\rho^{d}_{ij}}-s^{u}_{ij}e^{-i\rho^{u}_{ij}}$, which represents the analytic behavior of the CKM phase. The uncertainty in the relation is of order $O(\lambda^{2})\sim 4\%$, comparable to the current experimental precision. Comparisons with experimental data suggest that the hypothesis of some CP phases being maximal.
We also discussed relationships between the phase $\delta$ and unitarity triangles. As a result, several relations between the angles $\alpha,\beta,\gamma$ and $\delta$ are identified through other invariants $V_{il}V_{jm}V_{kn}/\det V_{\rm CKM}$. These general perturbative relations broadly cover phenomenological calculations, and therefore, the presented results have wide applicability for studies of CP violation in flavor physics and unified theories.